OpenStudy (anonymous):

The slope of the line containing the points (-2, 3) and (-3, 1) is

pooja195 (pooja195):

\[\huge~\rm~\frac{ x_2-x_1 }{ y_2-y_1 }=slope\]

OpenStudy (anonymous):

im honestly just trying to pass

pooja195 (pooja195):

Im honestly just trying to help

OpenStudy (anonymous):

just need answers

pooja195 (pooja195):

Just need you to learn.

OpenStudy (anonymous):

not really this is for summer school and i just need to pass

pooja195 (pooja195):

\[\huge~\rm~\frac{ -3+2 }{ 1-3}=slope\]

OpenStudy (anonymous):

-1/-2

pooja195 (pooja195):

Congrats you jusr did something by yourself and you learned.

OpenStudy (anonymous):

thats not even one of the options to choose from

pooja195 (pooja195):

What are the options?

OpenStudy (anonymous):

-2/5.. 1/2 ...2

pooja195 (pooja195):

-1/2 is right though :/